The Rander Company acquired a delivery truck for $21,164 on Jan 1. The truck has an estimated salvage value of $589 and an estimated useful life of 9 years. The company operates on a calendar year accounting period and uses double-declining-balance depreciation. What amounts of depreciation should be recorded for the first year?
Answer #1

Depreciation rate :

Straight line rate = 100/9 = 11.11%

Double decline rate = 11.11*2 = 22.22%

First year depreciation = 21164*22.22% = 4703